Mary went to a new shopping mall with a new $50.00 dollar bill. She bought two b-day cards for a total of $9.55 and a small OJ soda for .95 cents. She was required to pay an 8% state sales tax on the total purchase. What PERCENT of her hard earned money does she have left?
Thanks!

Answer by nerdybill(7384)

cost of items = 9.55 + .95 = 10.5
tax on items = 10.5(.08) = 0.84
.
Total amount spent = 0.84 + 10.5 = 11.34
.
Amount left = 50 - 11.34 = 38.66
.
Percent left = 38.66/50 * 100 = 0.7732 * 100 = 77.32%